Please help me if you can.
The area of a triangular sheet of paper is 28 square inches. One side of the triangle is 2 inches more than three times the lenght of the altitude to that side. Find the lenght of that side and the altitude to the side?
Thank you

Answer by scott8148(6628)

let x=altitude

"One side of the triangle is 2 inches more than three times the lenght of the altitude to that side"
__ 3x+2

area=1/2 times altitude times base __ 28=(1/2)(x)(3x+2)

dividing by 1/2 __ 56=3x^2+2x __ subtracting 56 __ 0=3x^2+2x-56

factoring __ 0=(3x+14)(x-4) __ x=-14/3 and x=4 __ negative value not realistic

altitude is 4 and side is 14